Week 49 Machine coordinators: J. Wenninger, B. Holzer (sick !)
 
  | 
   Date 
   | 
  
   Summary 
   | 
  
   Link to report from morning meeting 
   | 
 
 
  | 
   Thursday 8th Dec 
   | 
  
  
   - RP survey.
 
	- Xmas technical stop. 
	General access in all points except Pt 3 & 7.
 
   
   | 
  
   
	8:30 meeting 
    
   | 
 
 
  | 
   Wednesday 7th Dec 
   | 
  
  
   - 01:30 New attempt of 
	ramp with modified blow-up - failed again.
 
	- 03:00 STABLE BEAMS 
	#2350 Initial luminosity 4x1026 
	cm-2s-1 
 
	- 03:40 Beam dumped, 
	S81 FPA.
 
	- 08:40 STABLE BEAMS 
	#2351 Initial luminosity 4x1026 
	cm-2s-1 
 
	- 10:20 Beam dumped, 
	S45 trip.
 
	- Precycle. A few RF 
	trips in the 'shadow'.
 
	- 14:00 Refill for 
	physics.
 
	- 14:45 Beam dumped at 
	end of injection due to trip of RF module M1B2.
 
	- 15:00 Not worth to 
	refill - end of physics for 2012 !!!
 
	- Attempt to test the 
	combined ramp and squeeze: failed due to absence of beam. The ion beam is 
	lost between extraction of PS and injection into the SPS. Despite desperate 
	efforts, it was not possible to restore the ion beam in the SPS.
 
	- 18:00 End of LHC run 
	2011.
 
	- 18:30 Switching off 
	all circuits. LHC to access mode.
 
   
   | 
  
   
	8:30 meeting 
   | 
 
 
  | 
   Tuesday 6th Dec 
   | 
  
  
   - 06:00 Dump fill 2344.
 
	- QPS problem on 
	B10.L2. Reset OK. Precycle.
 
	- Quench test with 
	ions.
 
	- 3 attempts - no 
	quench !
 
	- 16:00 Preparing for 
	physics fill.
 
	- ADT gating test for 
	BBQ - not successful.
 
	- 18:00 Ramp lost due 
	to problem with longitudinal blowup settings.
 
	- 20:00 New ramp, again 
	new blow-up failed, bunches became way too long at the start - problem with 
	FB on bunch length.
 
	- 21:00 STABLE BEAMS 
	#2349 Initial luminosity 4x1026 
	cm-2s-1 
 
	- 22:40 Beam lost, trip 
	of RF module M2B1.
 
   
   | 
  
   
	8:30
  meeting 
   | 
 
 
  | 
   Monday 5th Dec 
   | 
  
  
   - 03:15 STABLE BEAMS 
	#2343 Initial luminosity 5x1026 
	cm-2s-1 
 
	- 09:30 Dump fill 2343.Integrated 
	luminosity ~6 ub-1.
 
	- Ion quench test 
	preparation.
 
	- 11:00 Beam scraping 
	at injection for CMS.
 
	- 12:00 Electrical 
	glitch due to fault on 18kV circuit breaker. Lost 4 sectors, RF, MKIs, 
	ATLAS. Collimators in strange state.
 
	- 12:30 Second glitch 
	when trying to switch back on. Lost cryo Pt18 as a consequence of the second 
	glitch.
 
	- Waiting for cryo 
	recovery.
 
	- 18:00 Cryo Pt18 
	recovered well ahead of initial prediction (24:00).
 
	- Preparing for 
	physics.
 
	- 20:00 Injecting.
 
	- 23:00 STABLE BEAMS 
	#2344 Initial luminosity 4.5x1026 
	cm-2s-1
 
   
   | 
  
   
	8:30 meeting 
   |